OpenStudy (anonymous):

I need help with 2 quick questions (questions are in comments)

OpenStudy (bookworm14):

the formula for circumference is \[2\pi r\] or 2 *3.14 * radius your radius is half of the diameter so for this problem #1 you radius is 14 (half of 28) now insert that into the equation 2 * 3.14 * 14 = 87.9 or if you round 88 the answer is A same strategy for #2 2 * 3.14 * 12.5 = 78.5 or (its already rounded) then answer for #2 is C Hope this helps! ~Bree :)

OpenStudy (bookworm14):

@flowers4547